How many stairs can a 63-kg person climb up with the energy contained in a candy bar (Ebar= 200 kcal)? Assume the height of each stair is 0.15 m and the efficiency of the person is 20%. Note: 1 kcal = 4184 J
Answer #1

Given Energy = 200 kcal = 200 * 4184 J = 836800 J

Now Energy Available to Man = 0.2 x 836800 = 167360 J

Now Let there are total n stairs .

So P.E of Man = Energy Available to Man

=> mg*n*h = 167360

=> n = (167360 / 63*9.81*0.15)

=> n = 1805 stairs -------- Answer